Changes listing for PSTricks

Version 0.94-Beta April 16, 1993 Timothy Van Zandt tvz@Princeton.EDU

The PSTricks package is still preliminary. The features list and user interface are unlikely to change much in the short term, but the internal code is not stable.

1 VERSION 0.94 (Beta)

1.1 Incompatible changes

1. \Rnode uses the graphics parameters href and vref to set the reference point, rather than having a coordinate argument.

1.2 New features

1. New graphics objects: \pstriangle, \psdiamond, \psdot.
2. New box frames: \pstribox, \psdiabox.
3. New nodes: \dotnode, \dianode, trinode.

4. Tree macros!

5. \psmatrix: Good for math diagrams and graphs.

6. Negative nodesep and offset measure distance from center of node.

7. New node connections: \ncbox, \ncarcbox.

8. New node label puts: \ncput, \naput, \nbput. (These replace \lput, \mput, \aput, \Aput, \bput and \Bput, which are still supported.)

9. New node label puts (especially good for trees): \thput, \tvput,

  \tlput, \trput, \taput, \tbput. (Make \ncline obsolete, but this is

still supported.)

10. \nput, for attaching labels to nodes.

11. New arrow styles: |<->| and |<*->|*.

12. New dot styles: diamond, diamond*, oplus, ominus, x, asterisk.

1.3 Obsolete but supported features

1. \lput, \mput, \aput, \Aput, \bput and \Bput.
2. \ncline.

1.4 Bug fixes

1. Fixed \ncangle so that it would work properly with non-right angle

  values of angleA and angleB.

1.5 Other changes

1. \OldPsput and \NewPsput are back in pstricks.tex.

2. The num in arrowsize=dim num, tbarsize=dim num, and dotsize=dim num is optional (default=0).